# TVR: Automotive System Requirement Traceability Validation and Recovery Through Retrieval-Augmented Generation

### v1

@@ 4.4.2. Traceability Recovery
To evaluate the performance of TVR in autonomous requirements traceability recovery, we manually verify their correctness (Formula 1). The evaluation metric is the proportion of correct traceability links, calculated as the number of links confirmed as correct through human verification divided by the total number of links identified by the model.
@@ 5.2. Best Prompt Strategy for Requirements Traceability Validation (RQ2)
Approach. This RQ explores the performance of different prompt engineering strategies for the requirement traceability validation task, specifically including CoT, few-shot, self-consistency, and RAG.
@@
Based on the results shown in Table 1, we select the best-performing model (i.e., Claude 3.5 Sonnet) for few-shot prompting and then evaluate self-consistency on this model. Specifically, we run the model ten times and apply a majority voting method, selecting the most frequent output among the ten results as the final output of the model (Wang et al., 2022).
@@
For few-shot prompting, the best-performing model is Claude 3.5 Sonnet. It outperforms all other models in terms of average performance across both categories. It achieves over 98% in precision, recall, and F1 score for the valid category, and over 86% for invalid category. However, applying the self-consistency method to Claude 3.5 Sonnet does not significantly improve its overall performance, as the F1 score improvement (in percentage) in both categories remains within 3%.
@@
5.3. Robustness to Unseen Requirements Variations (RQ3)
@@ 5.4. Performance on Requirements Traceability Recovery (RQ4)
Approach. As explained in Section 3.3, we employ a three-step pre-filtering process to reduce the number of requirement pairs we consider could have a missing link. For the remaining pairs, TVR predicts whether there is a valid traceability link between them. We then calculated the Correctness of pairs predicted as ‘Yes’.
Results. The three stages of pre-filtering yielded, in turn, 30,598, 14,494, and 1,919 requirement pairs. This demonstrates the effectiveness of our three-step pre-filtering, which significantly reduces infeasible pairs. Then 502 pairs out of 1,919 pairs are predicted to have valid traceability links. After manually verifying all 502 pairs, we obtained a Correctness of 85.50%, suggesting that most retrieved links are correct, with a limited number of false positives.
@@ 7. Threats to Validity
Internal Validity. The dataset was annotated by co-authors under the guidance of system engineers. However, manual annotation is inherently prone to errors, such as misinterpretation of requirements or inconsistencies in labeling. These annotation errors could affect the evaluation of TVR’s performance. Another internal threat arises from the model’s robustness evaluation, which is affected by the degree of variation among different categories. To mitigate this threat, we applied the concept of cross-validation to evaluate the model’s robustness across all four categories.
@@ 8. Related Work
Early automated traceability techniques relied on classical NLP methods, primarily leveraging information retrieval techniques (Guo et al., 2024). These methods establish potential traceability links by computing textual similarity between artifacts using models such as the vector space model (Antoniol et al., 2002; Mahmoud, 2015), latent semantic indexing (LSI) (Marcus and Maletic, 2003), latent Dirichlet allocation (LDA) (Asuncion et al., 2010), and hybrid approaches (Gethers et al., 2011; Moran et al., 2020). However, these traditional models struggle to capture deep semantics. Recent approaches have addressed this limitation by incorporating word embeddings and deep neural networks (Guo et al., 2017; Wang et al., 2018; Zhao et al., 2017; Hey et al., 2021). Additionally, some studies have explored active learning (Mills et al., 2019) and self-attention mechanisms (Zhang et al., 2021) for improving traceability link recovery.
@@
With the advent of LLMs such as GPT (OpenAI, 2023a, b), Llama (https://www.llama.com/, 2023), and Claude (https://claude.ai/, 2023), new opportunities have emerged for automating traceability while addressing the limitations of previous methods. Hey et al. (Hey et al., 2021) proposed the FTLR approach, which integrates word embedding similarity and Word Mover’s Distance (Kusner et al., 2015) to generate requirements-to-code trace links. Hey et al. (Hey et al., 2024) further enhances accuracy by incorporating a NoRBERT classifier to filter irrelevant requirement segments. Rodriguez et al. (Rodriguez et al., 2023) investigated the use of Claude, to directly perform traceability link recovery through natural language prompting. Their study demonstrated that generative models can not only recover trace links but also provide explanations for identified links. Fuchss et al. (Fuchß et al., 2025) recently introduced LiSSA, a RAG-enhanced LLM approach for traceability link recovery across requirements-to-code, documentation-to-code, and architecture documentation-to-models. LiSSA first retrieves relevant elements based on similarity values between embeddings, then generates LLM-based prompts to assess traceability links.
@@
In contrast to prior studies, our research focuses on validating traceability links as well as recovering missing links, addressing specific industry needs in the automotive sector: 1) Industry practitioners require a mechanism to verify the correctness of traceability links between stakeholder requirements and system requirements established by system engineers. 2) Both stakeholder and system requirements exhibit variations in the way they are expressed, with the possibility of encountering unseen variations in the future. 3) Both stakeholder and system requirements follow loose templates, with small differences in the way message and signal values are expressed, thus rendering information retrieval-based approaches based on term frequency and word overlapping ineffective. To address these challenges, we frame the validation task as a binary classification problem, where LLMs determine the validity of existing traceability links. LLMs are well-suited for this task due to their ability to understand and capture the subtle differences in messages and signals, in the presence of values among variations in requirements templates while simultaneously supporting explainability for traceability link validation.

### v2

@@ 4.4.2. Traceability Recovery
To evaluate TVR’s performance on automotive requirements traceability recovery, two of our authors manually verified its Correctness (Formula 5), and discussions were held to reach consensus when there was any disagreement. The evaluation metric is the proportion of correct traceability links, calculated as the number of links confirmed as correct through human verification divided by the total number of links identified by the model.
@@
4.5. Experiment Environment
All experiments were conducted on a laptop provided by our industry partner, running Windows 10, equipped with an Intel Core i7-11850H CPU at 2.5 GHz and 32 GB of RAM. Due to the data‐privacy concerns, we were limited to this single machine, which constrained our choice of text‐representation techniques. A temperature of 0 was applied to all LLMs to ensure consistent and reproducible outputs throughout the experiments.
@@ 5.2. Best Prompt Strategy for Requirements Traceability Validation (RQ2)
Approach. This RQ explores the performance of different prompt engineering strategies for the requirements traceability validation task, specifically including CoT, few-shot, self-consistency, and RAG.
@@
Based on the results shown in Table 2, we select the best-performing model (i.e., Claude 3.5 Sonnet) for few-shot prompting and then evaluate self-consistency on this model. Specifically, we run the model ten times and apply a majority voting approach, selecting the most frequent output among the ten results as the final output of the model (Wang et al., 2022).
@@
For few-shot prompting, the best-performing model is Claude 3.5 Sonnet. It outperforms all other models in terms of average performance across both categories. It achieves over 98% in precision, recall, and F1 score for the valid category, and over 86% for invalid category. However, applying the self-consistency approach to Claude 3.5 Sonnet does not significantly improve its overall performance, as the F1 score improvement (in percentage) in both categories remains within 3%.
@@
5.3. Robustness to Unseen Requirement Variations (RQ3)
@@ 5.4. Performance on Requirements Traceability Recovery (RQ4)
Approach. As explained in Section 3.3, we employ a three-step preprocessing process to reduce the number of requirement pairs we consider could have a missing link. For the remaining pairs, TVR predicts whether there is a valid traceability link between them. We then calculated the Correctness of pairs predicted as “Yes”.
Results. The three stages of preprocessing yielded, in turn, 30,598, 14,494, and 1,919 requirement pairs. This demonstrates the effectiveness of our three-step preprocessing, which significantly reduces infeasible pairs. Then 502 pairs out of 1,919 pairs are predicted to have valid traceability links. After manually verifying all 502 pairs, we obtained a Correctness of 85.50%, suggesting that most retrieved links are correct, with a limited number of false positives.
@@ 7. Threats to Validity
Internal Validity. The dataset was annotated by co-authors under the guidance of system engineers. However, manual annotation is inherently prone to errors, such as misinterpretation of requirements or inconsistencies in labeling. These annotation errors could affect the evaluation of TVR’s performance. To mitigate this threat, two of our authors independently labeled the dataset and resolved any conflicts through discussion. LLMs may produce different outputs for the same input due to their stochastic nature. We mitigated this threat by setting the temperature to 0 to ensure consistent and deterministic outputs. Another internal threat arises from the model’s robustness evaluation, which is affected by the degree of variation among different categories. To mitigate this threat, we applied the concept of cross-validation to evaluate the model’s robustness across all four categories.
@@ 8. Related Work
Early automated traceability techniques relied on classical NLP approaches, primarily leveraging information retrieval techniques (Guo et al., 2024). These approaches establish potential traceability links by computing textual similarity between artifacts using models such as the vector space model (Antoniol et al., 2002; Mahmoud, 2015), latent semantic indexing (LSI) (Marcus and Maletic, 2003), latent Dirichlet allocation (LDA) (Asuncion et al., 2010), and hybrid approaches (Gethers et al., 2011; Moran et al., 2020). However, these traditional models struggle to capture deep semantics. Recent approaches have addressed this limitation by incorporating word embeddings and deep neural networks (Guo et al., 2017; Wang et al., 2018; Zhao et al., 2017; Hey et al., 2021). Additionally, some studies have explored active learning (Mills et al., 2019) and self-attention mechanisms (Zhang et al., 2021) for improving traceability link recovery.
@@
With the advent of LLMs such as GPT (OpenAI, 2023a, b), Llama (https://www.llama.com/, 2023), and Claude (https://claude.ai/, 2023), new opportunities have emerged for automating traceability while addressing the limitations of previous approaches. Hey et al. (Hey et al., 2021) proposed the FTLR approach, which integrates word embedding similarity and Word Mover’s Distance (Kusner et al., 2015) to generate requirements-to-code trace links. Hey et al. (Hey et al., 2024) further enhances accuracy by incorporating a NoRBERT classifier to filter irrelevant requirement segments. Rodriguez et al. (Rodriguez et al., 2023) investigated the use of Claude, to directly perform traceability link recovery through natural language prompting. Their study demonstrated that generative models can not only recover trace links but also provide explanations for identified links. Fuchss et al. (Fuchß et al., 2025) recently introduced LiSSA, a RAG-enhanced LLM approach for traceability link recovery across requirements-to-code, documentation-to-code, and architecture documentation-to-models. LiSSA first retrieves relevant elements based on similarity values between embeddings, then generates LLM-based prompts to assess traceability links.
Similarly, in a recent study, Hey et al. (Hey et al., 2025) use GPT-4o to automate inter-requirement traceability link recovery (TLR) between high-level and low-level requirements. Their approach first retrieves similar requirements, then employs zero-shot and chain-of-thought prompting to verify the traceability links. However, existing approaches achieve limited accuracy, as they rely solely on simple zero-shot or chain-of-thought prompts combined with generic queries (e.g., “Is there a traceability link between…”). These approaches are inadequate for precisely verifying traceability links between DTC system and stakeholder requirements, as evidenced by the results in Table 2. DTC requirements often vary slightly in the character strings of messages and signals. Consequently, such general-purpose approaches struggle to capture the subtle nuances of domain-specific requirements.
@@
In contrast to prior studies, our research focuses on validating traceability links as well as recovering missing links, addressing specific industry needs in the automotive sector: 1) Industry practitioners require a mechanism to verify the correctness of traceability links between stakeholder requirements and system requirements established by system engineers. 2) Both stakeholder and system requirements exhibit variations in the way they are expressed, with the possibility of encountering unseen variations in the future. 3) Both stakeholder and system requirements follow loose templates, with small differences in the way message and signal values are expressed, thus rendering information retrieval-based approaches based on term frequency and word overlapping ineffective. To address these challenges, we frame the validation task as a binary classification problem, where LLMs determine the validity of existing traceability links. Moreover, we enhance the LLMs’ ability to understand the links between requirements by retrieving similar requirement pairs as in-context examples, which significantly improves performance compared to zero-shot and chain-of-thought prompting. LLMs are well-suited for this task due to their ability to understand and capture the subtle differences in messages and signals, in the presence of values among variations in requirements templates while simultaneously supporting explainability for traceability link validation.